# POLAR: A Pessimistic Model-based Policy Learning Algorithm for Dynamic Treatment Regimes

### v1

@@ 5.2 Simulation results
We first examine the performance of POLAR by examining the policy value derived from the estimated optimal policy. When the behavior policy $\pi_{b}$ is defined with $p=0.75$ and the offline data size is $n=200$, Figure 1 presents the policy value versus the number of iterations for POLAR under different values of the hyperparameter $c$ and all comparison methods, averaged over 100 repeated simulations. As the iteration increases, the policy values of POLAR with different choices of $c$ all converge. When no uncertainty penalization is applied ($c=0$), POLAR underperforms, likely due to overfitting in state-action regions poorly covered by the offline data. On the other extreme, with a large penalization parameter ($c=100$), the policy becomes overly conservative, prioritizing uncertainty minimization at the cost of reward. The best performance is achieved when $c=50$, where POLAR effectively balances between reward maximization and uncertainty penalization. This moderate level of pessimism helps the algorithm mitigate the effects of distributional shift while still learning a near-optimal policy. Compared to the baseline methods, POLAR consistently achieves higher policy values than DTR-Q, DDQN, and DTRreg, highlighting the benefits of incorporating a principled pessimism strategy and leveraging full historical information in offline RL.
Figure 1: Policy value vs. iterations for different values of $c$ ($p$=0.75, $n$=200).
@@
We then systematically analyze the performance of POLAR with respect to three factors: the parameter $p$, which determines the behavior policy $\pi_{b}$; the hyperparameter $c$, which controls the degree of pessimism; and the offline sample size $n$. For each combination of $(p,n,c)$, the average policy value across 100 repeated simulations is reported in Figure 2. When $p=0.95$ (Figure 2a), the behavior policy predominantly selects the optimal action, resulting in poor coverage of the state-action space and substantial distributional shift. In this setting, incorporating pessimism is critical: POLAR with moderate values of $c$ (e.g., $c=10$) outperforms both the no-penalty case ($c=0$) and the overly conservative setting ($c=100$). Conversely, when $p=0.55$, the behavior policy explores both actions more evenly, resulting in better state-action coverage. In this setting, uncertainty penalization is less critical, so smaller values of $c$ (e.g., $c=0,5,10$) yield the best performance, while larger values of $c$ (e.g., $c=50,100$) result in overly conservative policies and reduced performance, particularly when $n$ is small.
@@
Next, we observe that across all values of $p$ and $c$, POLAR’s performance improves as the offline sample size $n$ increases. When $n$ is sufficiently large, the policy values for different choices of $c$ begin to converge, indicating that the influence of the uncertainty penalty diminishes. This is because the uncertainty quantifier $\Gamma$ converges to zero at a rate of $\mathcal{O}_{p}(\sqrt{\log n/n})$, thereby reducing the impact of uncertainty penalization in large datasets. In contrast, when $n$ is small, POLAR significantly outperforms the baseline methods (DTR-Q, DTRreg, and DDQN), demonstrating its robustness in low-data regimes.
@@
Finally, we compare POLAR to the baseline methods in more detail. DDQN, which relies on the Markov assumption and does not incorporate historical information, shows relatively flat performance across all settings. This highlights its limitations in the DTR setup, where past trajectories contain critical information about patient history and treatment response that is essential for optimal decision-making. DTRreg approximates the Q-function using linear models, which limits its flexibility and results in consistently inferior performance. DTR-Q, on the other hand, improves steadily as $n$ increases and achieves performance comparable to POLAR when $n=100,000$ (not shown in the figure). This highlights a key distinction between model-based and model-free approaches: POLAR, by leveraging an explicit transition model and pessimism regularization, achieves near-optimal performance with far fewer samples, demonstrating the sample efficiency advantage of model-based approaches.
@@
These results demonstrate the importance of incorporating pessimism in offline DTR, particularly under limited trajectory coverage. When the behavior policy is skewed toward optimal actions (i.e., large $p$), offline data tend to underrepresent suboptimal regions, increasing the risk of overfitting. In such cases, a principled pessimism strategy helps mitigate distributional shift by regularizing uncertain estimates in poorly supported areas of the state-action space. This setting is especially relevant in medical applications, where observational data often reflect clinician-driven decisions that are near-optimal but selectively explored. By addressing this inherent coverage bias, POLAR offers a robust approach to policy learning from real-world clinical data.
@@ 6 Real Data Analysis
We applied the proposed POLAR method, using GP regression to estimate the transition dynamics $\widehat{P}_{k}(\cdot\mid h_{k},a_{k})$ and the corresponding uncertainty quantifiers $\Gamma_{k}(h_{k},a_{k})$ (Chang et al.,, 2021), across various values of the pessimism parameter $c\in\{0,5,10,20,50\}$. For comparison, we considered three baseline methods used in the simulation studies: DDQN, DTR-Q, and DTRreg. However, since DTRreg package cannot handle multiple action dimensions, we report results only for DDQN and DTR-Q. To assess model performance, we randomly split the dataset into 50% training and 50% testing sets. The training set was used to train the algorithms, while the testing set was employed to evaluate the performance of the learned policies through off-policy evaluation (OPE) with importance sampling (Uehara et al.,, 2022). This procedure was repeated $N=200$ times with different random splits, and the final results are reported as averages across these repetitions.
@@
Figure 3: Policy value versus number of iterations for POLAR with different values of $c$, compared to DDQN and DTR-Q. All policy values are evaluated using OPE.
@@
Figure 4: The action distributions generated by the physician policy, POLAR ($c=20$), DTR-$Q$, and DDQN across three decision stages.
@@
Figure 3 presents the policy value versus iteration $t$ for the proposed POLAR with different values of $c$, alongside two baseline methods, DTR-Q and DDQN. Overall, POLAR achieves higher policy values than both baselines, demonstrating its effectiveness in real-world offline settings. Introducing a moderate pessimistic penalty generally improves performance: configurations with $c=5,10$, and 20 outperform the no-penalty case ($c=0$) throughout the iterations. Among these, $c=20$ attains the highest policy value, suggesting that incorporating a conservative bias can be beneficial in managing uncertainty. However, excessively large penalties (e.g., $c=50$) lead to degraded performance, demonstrating the importance of balancing uncertainty penalization with reward maximization in offline policy learning.
@@
To further compare the methods, we analyzed the action distributions generated by the physician policy (i.e., the behavior policy), POLAR ($c=20$), DTR-$Q$, and DDQN across three decision stages (Figure 4). The physician policy demonstrated distinct treatment patterns: lower vasopressor dosages with varying IV fluid levels in Stage 1, followed by a shift toward higher vasopressor dosages in later stages. These changes reflect clinicians’ dynamic responses to evolving patient conditions and the importance of history-dependent decision-making in sepsis management.
@@
POLAR captured key aspects of these temporal treatment shifts, adapting its action distributions across stages to reflect increasingly aggressive vasopressor use while maintaining diversity in fluid levels. By incorporating uncertainty-aware regularization and leveraging historical patient information, POLAR avoided overcommitting to uncertain actions and effectively balanced exploitation and exploration. Its distributions demonstrated clinically interpretable structure and were aligned with domain expectations for stage-specific intervention patterns. In contrast, DTR‐Q and DDQN exhibited limited adaptability across stages. DTR‐Q failed to learn effective physician actions in stage 1, producing a highly dispersed distribution and recommending large vasopressor doses that would be dangerously aggressive in practice. DDQN displayed nearly identical distributions across all three stages and did not learn the gradual vasopressor escalation trend, suggesting its weakness in handling short sequential decision processes. These findings demonstrate POLAR’s ability to model the dynamic, context‐aware nature of clinical decision‐making. By leveraging historical information, sampling unexplored state–action combinations, and penalizing uncertain actions, POLAR learns interpretable, adaptive treatment strategies that more closely align with real‐world clinical expertise.
@@ Acknowledgment
This work is supported in part by National Institute of Health grant R01MH128085.
@@ A.2.1 Model Specification and Estimation
Unlike the example in Section A.1, we no longer assume a parametric form for $P_{k}$. Instead, we consider a nonparametric setting. We assume the transition dynamic $P_{k}(s_{k+1}\mid h_{k},a_{k})$ is structured as: $s_{k+1}=g_{k}(h_{k},a_{k})+\varepsilon_{k}$, where $\varepsilon_{k}\sim N(0,\sigma^{2}I)$.
@@
Under the GP prior, the posterior distribution over the unknown function $g$ remains a Gaussian process with posterior mean function $\widehat{g}(\cdot)$ and covariance function $\widehat{h}(\cdot,\cdot)$, given by:
@@ A.2.2 Theoretical Properties
3 Construct the modified DTR model $\widetilde{M}=(\widehat{P},\widetilde{r})$, with $\widetilde{r}_{k}=\widehat{r}_{k}-\widetilde{c}_{k}\Gamma_{k}$
@@ C.1.1 Proofs of Theorem 1
Suppose Assumption 1 holds. Define $\widetilde{c}_{k}:=\sum_{j=k}^{K}\|\overline{r}_{j}\|_{\infty}$, and let $\widetilde{r}_{k}:=\widehat{r}_{k}-\widetilde{c}_{k}\Gamma_{k}$. Consider the modified model $\widetilde{M}=(\widehat{P},\widetilde{r})$. Then, for any policy $\pi$, the following inequality holds with probability at least $1-K\delta$:
@@
    $V_{k}^{\pi}(h_{k})-\widehat{V}_{k}^{\pi}(h_{k})=\sum_{t=k}^{K-1}\mathbb{E}_{(h_{t},a_{t})\sim d_{t,P}^{\pi}(\cdot,\cdot\mid H_{k}=h_{k})}\left[\mathbb{E}_{h^{\prime}\sim P_{t}(\cdot\mid h_{t},a_{t})}\left[\widehat{V}_{t+1}^{\pi}(h^{\prime})\right]-\mathbb{E}_{h^{\prime}\sim\widehat{P}_{t}(\cdot\mid h_{t},a_{t})}\left[\widehat{V}_{t+1}^{\pi}(h^{\prime})\right]\right].$
@@
Suppose Assumption 2 holds, and let $\widetilde{M}=(\widehat{P},\widetilde{r})$ be the modified DTR model constructed in Algorithm 1. Assume that the step sizes are constant over iterations, i.e., $\eta_{k}^{(t)}=\eta_{k}$ for all $t$. Then, Algorithm 1 returns a policy $\pi^{(T)}$ that satisfies the following performance guarantee:
@@ C.1.2 Proof of Proposition 1
[Proposition 1 Restated] Suppose $\widetilde{M}=(\widehat{P},\widetilde{r})$ is the modified DTR model constructed in Algorithm 2, and that Assumptions S2S3S4S5, and S6 hold. Let $\eta_{k}^{(t)}=\frac{\sqrt{\log|\mathcal{A}_{k}|}}{\widetilde{q}_{k}\sqrt{T}}$, constant in $t$. If $L_{k}\asymp\left(\frac{m_{k}}{\log m_{k}}\right)^{\overline{d}_{k}/(2p+\overline{d}_{k})}$ and $p\delta_{k}\geq\overline{d}_{k}$, then Algorithm 2 returns a policy $\pi^{(T)}$ satisfying:
@@
C.2 Proof of Theorem 2 in Section 4.2.1
@@
There exist constants $c_{1}$, $c_{2}$ such that the following two conditions hold for all $1\leq k\leq K$: (i) $dim(\mathcal{S}_{k})\leq c_{1}$; (ii) $d_{k}^{(s)}\leq c_{2}k$.
Assumption S7 (i) trivially holds because $K$ is finite. Therefore, $dim(\mathcal{S}_{k})=\mathcal{O}(1)$, which further implies $dim(\mathcal{H}_{k}\times\mathcal{A}_{k})=k+\sum_{j=1}^{k}dim(\mathcal{S}_{j})=\mathcal{O}(k)$. Recall that $\phi_{k}:\mathcal{H}_{k}\times\mathcal{A}_{k}\to\mathbb{R}^{d_{k}^{(s)}}$, so Assumption S7 (ii) is a mild condition that naturally requires the feature mapping $\phi_{k}$ to reserve the dimensional structure of $\mathcal{H}_{k}\times\mathcal{A}_{k}$.
@@
The feature map $\phi_{k}:\mathcal{H}_{k}\times\mathcal{A}_{k}\to\mathbb{R}^{d_{k}^{(s)}}$ is uniformly bounded. There exists a constant $R>0$ such that
$\|\phi_{k}(h,a)\|_{2}^{2}\;\leq\;R^{2}d_{k}^{(s)},\quad\text{for all }(h,a)\in\mathcal{H}_{k}\times\mathcal{A}_{k}.$
@@
Since $\Lambda_{k,n}$ is the regularized empirical covariance matrix based on $n$ i.i.d. samples, we have $\det(\Lambda_{k,n})=\mathcal{O}(n^{d_{k}^{(s)}})$. Therefore, under Assumption S7 (ii), it follows that $\beta_{k,n}\lesssim\sqrt{d_{k}^{(s)}\log n}\lesssim\sqrt{k\log n}$. Therefore,
$\Gamma_{k}(h_{k},a_{k})\lesssim\sqrt{k\log n}\cdot\sqrt{\phi_{k}(h_{k},a_{k})^{\top}\Lambda_{k,n}^{-1}\phi_{k}(h_{k},a_{k})}.$ (S29)
@@
$\mathbb{E}_{(h_{k},a_{k})\sim d_{P^{*},k}^{\pi^{\dagger}}}\Gamma_{k}(h_{k},a_{k})\lesssim\sqrt{k\log n\cdot\mathbb{E}_{(h_{k},a_{k})\sim d_{P^{*},k}^{\pi^{\dagger}}}\phi_{k}(h_{k},a_{k})^{\top}\Lambda_{k,n}^{-1}\phi_{k}(h_{k},a_{k})}.$ (S30)
@@
$\mathbb{E}_{(h_{k},a_{k})\sim d_{P^{*},k}^{\pi^{\dagger}}}\Gamma_{k}(h_{k},a_{k})\lesssim k\sqrt{\frac{C_{k}\log n}{n}}.$ (S31)
@@
$\|\Gamma_{k}\|_{\infty}\lesssim k\sqrt{\frac{\log n}{n}}.$ (S32)
@@
$b_{k}=2\widetilde{c}_{k}+\sum_{j=k+1}^{K}\widetilde{c}_{j}\|\Gamma_{j}\|_{\infty}=\mathcal{O}\left(K+\frac{K^{3}\sqrt{\log n}}{\sqrt{n}}\right).$
@@
$\left(V_{M^{*}}^{\pi^{\dagger}}-V_{\widetilde{M}}^{\pi^{\dagger}}\right)+\left(V_{\widetilde{M}}^{\pi}-V_{M^{*}}^{\pi}\right)\leq{\mathcal{O}}\left(\frac{K^{3}\sqrt{\log n}}{\sqrt{n}}+\frac{K^{5}}{n}\right),$
which completes the proof. ∎
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}\widetilde{\mathcal{O}}\left(\frac{K^{3}}{\sqrt{n}}+\frac{K^{5}}{n}\right)+\mathcal{O}\left(\frac{K}{\sqrt{T}}\right)+K\sqrt{T}\cdot\mathcal{O}_{p}\left(\left(\frac{\log m_{k}}{m_{k}}\right)^{\alpha_{1}}\right),$

### v2

@@ 5.2 Simulation results
We first examine the performance of POLAR by examining the policy value derived from the estimated optimal policy. When the behavior policy $\pi_{b}$ is defined with $p=0.75$ and the offline data size is $n=200$, Figure 1 presents the policy value versus the number of iterations for POLAR under different values of the hyperparameter $c$ and all comparison methods, averaged over 100 repeated simulations. As the iteration increases, the policy values of POLAR with different choices of $c$ all converge. When no uncertainty penalization is applied ($c=0$), POLAR underperforms, likely due to overfitting in state-action regions poorly covered by the offline data. On the other extreme, with a large penalization parameter ($c=100$), the policy becomes overly conservative, prioritizing uncertainty minimization at the cost of reward. The best performance is achieved when $c=50$, where POLAR effectively balances between reward maximization and uncertainty penalization. This moderate level of pessimism helps the algorithm mitigate the effects of distributional shift while still learning a near-optimal policy. Compared to the baseline methods, POLAR consistently achieves higher policy values than DTR-$Q$, DDQN, DTRreg, MOPO, and MILO, highlighting the benefits of incorporating a principled pessimism strategy and leveraging full historical information in offline RL.
Figure 1: Policy value vs. iterations for different values of $c$ (pp=0.75, nn=200).
@@
We then systematically analyze the performance of POLAR with respect to three factors: the parameter $p$, which determines the behavior policy $\pi_{b}$; pessimism hyperparameter $c$, which controls the degree of pessimism in the reward function; and the offline sample size $n$. For each combination of $(p,n,c)$, the average policy value across 100 repeated simulations is reported in Figure 2. When $p=0.95$ (Figure 2a), the behavior policy selects the optimal action with high probability, leading to poor state-action coverage and substantial distributional shift. In this setting, incorporating pessimism is important: POLAR with a non-zero penalty ($c>0$) outperforms the unpenalized case ($c=0$) across all sample sizes $n$. In contrast, when $p=0.55$, the behavior policy explores the actions more evenly, providing better state-action coverage. In this setting, uncertainty penalization becomes less critical. Small penalties (e.g., $c\in\{0,5,10\}$) achieve the best performance, while larger penalties (e.g., $c\in\{50,100\}$) lead to overly conservative policies and reduced performance, especially at small sample sizes.
@@
Next, we observe that across all values of $p$ and $c$, POLAR’s performance improves as the offline sample size $n$ increases. When $n$ is sufficiently large, the policy values for different choices of $c$ begin to converge, indicating that the influence of the uncertainty penalty diminishes. This is because the uncertainty quantifier $\Gamma$ converges to zero at a rate of $\mathcal{O}_{p}(\sqrt{\log n/n})$, thereby reducing the impact of uncertainty penalization in large datasets. In contrast, when $n$ is small, POLAR significantly outperforms the baseline methods (DTR-$Q$, DTRreg, DDQN, MOPO and MILO), demonstrating its robustness in low-data regimes.
@@
We further compare POLAR to the baseline methods in more detail. DDQN, which relies on the Markov assumption and does not incorporate historical information, shows relatively flat performance across all settings. This highlights its limitations in the DTR setup, where past trajectories contain critical information about patient history and treatment response that is essential for optimal decision-making. DTRreg approximates the Q-function using linear models, which limits its flexibility and results in consistently inferior performance. DTR-$Q$, on the other hand, improves steadily as $n$ increases and achieves performance comparable to POLAR when $n=100,000$ (not shown in the figure). This highlights a key distinction between model-based and model-free approaches: POLAR, by leveraging an explicit transition model and pessimism regularization, achieves near-optimal performance with far fewer samples, demonstrating the sample efficiency advantage of model-based approaches. While the policy values of MOPO and MILO increase with sample size $n$, they remain consistently lower than those estimated by POLAR. This result is expected. MILO focuses on imitating the behavior policy, which can constrain performance improvement. MOPO, designed for stationary Markov decision processes, relies on a single time-homogeneous transition model, introducing model mismatch in the multi-stage DTR setting. In contrast, POLAR is explicitly designed for finite-horizon, history-dependent DTRs, modeling stage-specific transition kernels ${P_{k}(h_{k+1}\mid h_{k},a_{k})}_{k=1}^{K}$.
@@
These results demonstrate the importance of incorporating pessimism in DTR, particularly under limited trajectory coverage. When the behavior policy is skewed toward optimal actions (i.e., large $p$), offline data tend to underrepresent suboptimal regions, increasing the risk of overfitting. In such cases, a principled pessimism strategy helps mitigate distributional shift by regularizing uncertain estimates in poorly supported areas of the state-action space. This setting is especially relevant in medical applications, where observational data often reflect clinician-driven decisions that are near-optimal but selectively explored.
Finally, we evaluate POLAR’s robustness through a sensitivity analysis. While the simulated data follow a linear transition model, we intentionally apply POLAR using a GP to model transitions, creating a model mismatch. Under this misspecification, POLAR continues to outperform all alternative methods, confirming its effectiveness. More details are provided in Supplement F.3.
@@ 6 Real Data Analysis
We applied the proposed POLAR method, using GP regression to estimate the transition dynamics $\widehat{P}_{k}(\cdot\mid h_{k},a_{k})$ and the corresponding uncertainty quantifiers $\Gamma_{k}(h_{k},a_{k})$ (Chang et al.,, 2021), across various values of the pessimism parameter $c\in\{0,5,10,20,50\}$. For comparison, we considered five baseline methods used in the simulation studies: DDQN, DTR-$Q$, DTRreg, MOPO and MILO. However, since DTRreg package cannot handle multiple action dimensions, we report results only for DDQN, DTR-$Q$, MOPO, and MILO. To assess model performance, we randomly split the dataset into 50% training and 50% testing sets. The training set was used to train the algorithms, while the testing set was employed to evaluate the performance of the learned policies through off-policy evaluation (OPE) with importance sampling (Uehara et al.,, 2022). This procedure was repeated $N=200$ times with different random splits, and the final results are reported as averages across these repetitions.
@@
Figure 3: Policy value versus number of iterations for POLAR with different values of $c$, compared to DDQN, DTR-$Q$, MOPO, and MILO. All policy values are evaluated using OPE.
@@
Figure 4: The action distributions generated by the physician policy, POLAR ($c=20$), DTR-$Q$, DDQN, MILO and MOPO across three decision stages.
@@
Figure 3 presents the policy value versus iteration $t$ for the proposed POLAR with different values of $c$, alongside four methods, DDQN, DTR-$Q$, MOPO, and MILO. Overall, POLAR achieves higher policy values than all baselines, demonstrating its effectiveness in real-world offline settings. Introducing a moderate pessimistic penalty generally improves performance: configurations with $c=5,10$, and 20 outperform the no-penalty case ($c=0$) throughout the iterations.
Among these, $c=20$ attains the highest policy value, suggesting that incorporating a conservative bias can be beneficial in managing uncertainty. However, excessively large penalties (e.g., $c=50$) lead to degraded performance, demonstrating the importance of balancing uncertainty penalization with reward maximization in offline policy learning. In practice, we recommend running POLAR with a candidate set of $c$ values and selecting the optimal one via cross-validated grid search.
@@
To further compare the methods, we analyzed the action distributions generated by the physician policy (i.e., the behavior policy), POLAR ($c=20$), DDQN, DTR-$Q$, MOPO, and MILO across three decision stages (Figure 4). The physician policy demonstrated distinct treatment patterns: lower vasopressor dosages with varying IV fluid levels in Stage 1, followed by a shift toward higher vasopressor dosages in later stages. These changes reflect clinicians’ dynamic responses to evolving patient conditions and the importance of history-dependent decision-making in sepsis management.
@@
POLAR captured key aspects of these temporal treatment shifts, adapting its action distributions across stages to reflect increasingly aggressive vasopressor use while maintaining diversity in fluid levels. By incorporating uncertainty-aware regularization and leveraging historical patient information, POLAR avoided overcommitting to uncertain actions and effectively balanced exploitation and exploration. Its distributions demonstrated clinically interpretable structure and were aligned with domain expectations for stage-specific intervention patterns. In contrast, DTR‐Q, DDQN, MILO, and MOPO exhibit limited adaptability across decision stages. DTR‐Q fails to learn effective physician actions in Stage 1, producing a highly dispersed action distribution and recommending excessively large vasopressor doses that would be dangerously aggressive in practice. DDQN displays nearly identical action distributions across all three stages and fails to capture the gradual vasopressor escalation pattern, highlighting its weakness in short-horizon sequential decision problems. MILO produces action distributions closely resembling the clinician’s, but this imitation-driven design yields lower OPE values than POLAR. MOPO similarly generates nearly invariant action distributions across decision stages, indicating limited adaptability to the stage-specific dynamics of clinical DTRs. These findings demonstrate POLAR’s ability to model the dynamic, context‐aware nature of clinical decision‐making. By leveraging historical information, sampling unexplored state–action combinations, and penalizing uncertain actions, POLAR learns interpretable, adaptive treatment strategies that more closely align with real‐world clinical expertise.
@@ Acknowledgment
This work is supported in part by National Institute of Health grant R01MH128085 and R01AI197147.
@@ A.2.1 Model Specification and Estimation
Unlike the example in Section A.1, we no longer assume a parametric form for $P_{k}$. Instead, we consider a nonparametric setting. We assume the transition dynamic $P_{k}(s_{k+1}\mid h_{k},a_{k})$ is structured as: $s_{k+1}=g_{k}(h_{k},a_{k})+\varepsilon_{k}$, where $\varepsilon_{k}\sim\mathcal{N}(0,\sigma^{2}I)$.
@@
[…]
@@ A.2.2 Theoretical Properties
[…]
@@ C.1.1 Proofs of Theorem 1
[…]
@@
    $V_{k}^{\pi}(h_{k})-\widehat{V}_{k}^{\pi}(h_{k})$ $=\sum_{t=k}^{K-1}\mathbb{E}_{(h_{t},a_{t})\sim d_{t,P}^{\pi}(\cdot,\cdot\mid H_{k}=h_{k})}\Biggl[\mathbb{E}_{h^{\prime}\sim P_{t}(\cdot\mid h_{t},a_{t})}\bigl[\widehat{V}_{t+1}^{\pi}(h^{\prime})\bigr]$
    $\qquad-\mathbb{E}_{h^{\prime}\sim\widehat{P}_{t}(\cdot\mid h_{t},a_{t})}\bigl[\widehat{V}_{t+1}^{\pi}(h^{\prime})\bigr]\Biggr].$
@@
[…]
@@ C.1.2 Proof of Proposition 1
[…]
@@
C.2 Proof of Theorem 2 and Theorem 3 in Section 4.2.1
@@
There exist constants $c_{1}$, $c_{2}$ such that the following two conditions hold for all $1\leq k\leq K$: (i) $dim(\mathcal{S}_{k})\leq c_{1}$; (ii) $\dim(\phi_{k})\leq c_{2}k$.
Assumption S7 (i) trivially holds because $K$ is finite. Therefore, $dim(\mathcal{S}_{k})=\mathcal{O}(1)$, which further implies $dim(\mathcal{H}_{k}\times\mathcal{A}_{k})=k+\sum_{j=1}^{k}dim(\mathcal{S}_{j})=\mathcal{O}(k)$. Recall that $\phi_{k}:\mathcal{H}_{k}\times\mathcal{A}_{k}\to\mathbb{R}^{\dim(\phi_{k})}$, so Assumption S7 (ii) is a mild condition that naturally requires the feature mapping $\phi_{k}$ to reserve the dimensional structure of $\mathcal{H}_{k}\times\mathcal{A}_{k}$.
@@
The feature map $\phi_{k}:\mathcal{H}_{k}\times\mathcal{A}_{k}\to\mathbb{R}^{\dim(\phi_{k})}$ is uniformly bounded. There exists a constant $R>0$ such that
$\|\phi_{k}(h,a)\|_{2}^{2}\;\leq\;R^{2}\dim(\phi_{k}),\quad\text{for all }(h,a)\in\mathcal{H}_{k}\times\mathcal{A}_{k}.$
@@
Since $\Lambda_{k,n}$ is the regularized empirical covariance matrix based on $n$ i.i.d. samples, we have $\det(\Lambda_{k,n})=\mathcal{O}(n^{d_{k}^{(s)}})$.
Therefore, under Assumption S7 (ii), it follows that $\beta_{k,n}\lesssim\sqrt{d_{k}^{(s)}\log n}\lesssim\sqrt{\log n}$. Hence,
$\Gamma_{k}(h_{k},a_{k})\lesssim\sqrt{\log n}\cdot\sqrt{\phi_{k}(h_{k},a_{k})^{\top}\Lambda_{k,n}^{-1}\phi_{k}(h_{k},a_{k})}.$ (S29)
@@
$\mathbb{E}_{(h_{k},a_{k})\sim d_{P^{*},k}^{\pi^{\dagger}}}\Gamma_{k}(h_{k},a_{k})\lesssim\sqrt{\log n\cdot\mathbb{E}_{(h_{k},a_{k})\sim d_{P^{*},k}^{\pi^{\dagger}}}\phi_{k}(h_{k},a_{k})^{\top}\Lambda_{k,n}^{-1}\phi_{k}(h_{k},a_{k})}.$ (S30)
@@
$\mathbb{E}_{(h_{k},a_{k})\sim d_{P^{*},k}^{\pi^{\dagger}}}\Gamma_{k}(h_{k},a_{k})\lesssim\sqrt{\frac{C_{k}\log n}{n}}.$ (S31)
@@
$\|\Gamma_{k}\|_{\infty}\lesssim\sqrt{\frac{k\log n}{n}}.$ (S32)
@@
$b_{k}=2\widetilde{c}_{k}+\sum_{j=k+1}^{K}\widetilde{c}_{j}\|\Gamma_{j}\|_{\infty}=\mathcal{O}\left(K+\frac{K^{5/2}\sqrt{\log n}}{\sqrt{n}}\right).$
@@
$\left(V_{M^{*}}^{\pi^{\dagger}}-V_{\widetilde{M}}^{\pi^{\dagger}}\right)+\left(V_{\widetilde{M}}^{\pi}-V_{M^{*}}^{\pi}\right)\leq{\mathcal{O}}\left(\frac{K^{2}\sqrt{\log n}}{\sqrt{n}}+\frac{K^{7/2}\log n}{n}\right),$
which completes the proof.
@@
$\stackrel{{\scriptstyle(i)}}{{\leq}}\widetilde{\mathcal{O}}\!\left(\frac{K^{2}}{\sqrt{n}}+\frac{K^{7/2}}{n}\right)+\mathcal{O}\!\left(\frac{K}{\sqrt{T}}\right)+K\sqrt{T}\cdot\mathcal{O}_{p}\!\left(\left(\frac{\log m_{k}}{m_{k}}\right)^{\alpha_{1}}\right),$